OpenStudy (anonymous):

Please help.... Identify the standard form of the equation of a line which has a slope of -1/2 and goes through (3,2) a.)x+2y=7 b.)x-2y=7 c.)2x-y=7 d.)2x+y=4

OpenStudy (anonymous):

y- b =m( x-a)

OpenStudy (anonymous):

m is slope n line is passing through the point (a,b)

OpenStudy (anonymous):

y-2= -1/2(x-3)

OpenStudy (anonymous):

2y-4=-x+3 2y=-x+7

OpenStudy (anonymous):

x+2y=7

OpenStudy (anonymous):

y=-1/2x +C (constant) this means that it has slope of -1/2 to get the constant you only need to check what happens at x=3 at x=3 it is -1/2*3+C=2 so C is 7/2 multiplying with 2 give 2y=-x+7 So the answer is A.

OpenStudy (anonymous):

thank you!